# JEL: C81, O32

### v3

@@
We develop a method for assigning high-quality labels to unstructured text. The method is based on fine-tuning an open-source language model with data extracted from a proprietary large language model. We apply this method to construct a census of published clinical trials. We revisit a literature that contends that pharmaceutical research productivity is declining, based on measured increases in the quantity of clinical trials, which outpace trends in output. In our data, the quantity and composition of trials are stable since 2010. Previous measurements are an artifact of biases driven by shifts in the composition of other forms of research.
@@
Perhaps the most serious task facing empirical work in the area of “technological change” and “invention and innovation” is the construction and interpretation of measures (indices) of advances in knowledge. — Pakes and Griliches,, 1980
@@ 1. Introduction
Considerable progress in the measurement of innovation has been made in the half-century since the publication of Pakes and Griliches, (1980). This progress has been enabled in large part by the proliferation of large-scale corpora of patents and scientific publications. Yet measurement remains an active constraint for empirical research (Bryan and Williams,, 2021). Patents and publications are attractive proxies for innovation, as they necessarily relate to instances of invention and research. Their utility is limited, however, by heterogeneity in their technical content and value (see e.g., Pakes,, 1986; Griliches,, 1990, 1994). Consequently, trends in counts of documents are often confounded by trends in their composition (Lerner and Seru,, 2022). Correction of this confounding can be viewed as a classification problem: how accurately can unstructured text be mapped to specific quantities of interest? [^1] The most prominent example of such a classification effort is the construction of the National Bureau of Economic Research (NBER) Patent Data, which, among other things, maps patent records to technology classes (Hall et al.,, 2001). Earlier, Scherer, (1984) classified, by hand, over 15,000 patents according to their technological category.
@@
In this paper, we demonstrate that large language models can facilitate classification tasks of this kind, achieving an accuracy and precision comparable to hand-made labels at a research-relevant scale and budget. We develop a procedure for constructing task-specific large language models, particularly suited to data construction problems that involve classification. The class of methods that we consider has two steps. First, a comparatively small number of labels are extracted from a large, typically proprietary, language model. [^2] By proprietary large language model, we are referring, at present, to models such as OpenAI’s GPT 3.5 and GPT-4. Alternative models include Anthropic’s Claude 3.5 Opus, Google’s Gemini, among many others. Second, these labels are used as training data for a smaller, more efficient, open-source language model. [^3] Here, we are referring to classes of open-source language models such as RoBERTa, based on Google’s BERT model, among others. This process is an instance of model distillation, which is a growing area of methodological research in computer science (Taori et al.,, 2023; Xu et al.,, 2024). This procedure allows researchers to capture the benefits of complex, frontier models, while retaining the resource efficiency and transparency of lighter-weight open-source models.
@@
The primary contribution of this paper is the application of this procedure in a context that is emblematic of the econometric challenges faced by studies of innovation: measurement of trends in clinical trial production. We show that standard methods for data classification identify clinical trials in publication data imprecisely. These imprecise measurements indicate, spuriously, that the quantity of clinical trials has increased sharply since 2010. In our data, the quantity, quality, and composition of clinical trials have been stable year-on-year.
When considered alongside recent evidence on the stability of drug development costs during this period (Sertkaya et al.,, 2024), these new measurements of clinical trial production call into question the conclusions of an influential literature that finds diminishing returns to pharmaceutical research (Scannell et al.,, 2012; DiMasi et al.,, 2016; Bloom et al.,, 2020). [^4] Goldin et al., (2024) survey evidence related to declining productivity across the economy and note that, in the pharmaceutical sector, evidence for declining productivity primarily relies on measures of total numbers of clinical trials and estimates of R&D expenditures. The conclusions of this literature are based on measures that suggest that the quantity of clinical trials is increasing substantially, in a way unmatched by trends in measures of outputs. In light of our new data, we find that there is little rigorous empirical support for falling pharmaceutical research productivity in the most recent period. This finding is directly relevant to a set of active policy debates, which were initiated, in no small part, in response to concerns that the production of innovative output in this sector is slowing down. These debates include proposals that would restructure public subsidies administered by the National Institutes for Health (NIH) (e.g., Executive Office of the President,, 2019; Moss et al.,, 2020; The Economist,, 2022), relax the regulatory standards for approval of new medicines by the U.S. Food and Drug Administration (FDA) (e.g., Roy,, 2021; Chertman,, 2023; Graboyes and Reimschisel,, 2017), and prevent the Centers for Medicare and Medicaid Services (CMS) from negotiating drug prices (e.g., York,, 2023; Long,, 2023; Inflation Reduction Act,, 2022).
Clinical trials are an especially attractive measure of research in medicine (Bryan and Williams,, 2021). In the pharmaceutical industry, clinical trials are discrete, regulated experiments that generate useful statistical information about new and existing medicines (Chavez-MacGregor and Giordano,, 2016). Trials top the hierarchy of medical evidence employed by firms, regulators, physicians, and patients (Jones and Podolsky,, 2015). Since the early 2000s, clinical trials have been reported in a consistent, standardized format across scientific publications and, with less consistency, in administrative databases (DeVito et al.,, 2020; DeVito and Goldacre,, 2021). Trials have, thus, been widely used to measure relative and absolute differences in the composition of research investment (see Finkelstein, (2004) and Budish et al., (2015) for two influential examples). Studies of information production and disclosure have examined the substantive content of trials (see e.g., Alsan et al., (2024), Oostrom, (2024), and Kao and Oostrom, (2024)). Recently, and prominently, measures of clinical trial quantities have been used to support arguments that medical and pharmaceutical research productivity are declining over time (Scannell et al.,, 2012; Bloom et al.,, 2020).
Nevertheless, it is remarkably difficult to generate accurate measures of trends in clinical trial production. Administrative databases—such as ClinicalTrials.gov—have experienced large shifts in compliance with reporting regulations over time, which make it difficult to distinguish genuine increases in research effort from greater adherence to reporting requirements (see e.g., DeVito et al.,, 2020). Proprietary databases (e.g., Cortellis,, 2024) take as an input the contents of publicly-maintained registries and, thus, are confounded by the same shifts in compliance.
Given these limitations, scientific publication data have served as a natural alternative or complement in recent empirical studies (e.g., Lichtenberg,, 2018; Bloom et al.,, 2020; Kao,, 2024; Guzman et al.,, 2024). With unstructured publication text data, the challenge becomes, again, one of classification: of the millions of records in large-scale publication databases, which disclose the results of a clinical trial? Manually identifying clinical trial records at a scale sufficient to study aggregate trends is infeasible. [^5] A small number of papers curate select samples of clinical trial data, for specific research questions. For example, Oostrom, (2024) collects records of psychiatric clinical trials as identified in prominent meta-analyses for a study of the relationship between sources of clinical trial funding and reported drug efficacy. We show that standard methods—based on publication metadata and machine learning tools—perform poorly. In particular, we show that keyword-based searches, including methods employed by Bloom et al., (2020), are very imprecise and generate spurious measurements of increases in the quantity of clinical trials. [^6] Specifically, Bloom et al., (2020) categorize many records as clinical trials that discuss trials (e.g., meta-analyses and literature reviews). We find that changes in the scientific publishing landscape, discussed in Section 4, increase the size of this categorization error over time. Specialized classifiers developed for use in the health services literature, as in Thomas et al., (2021), similarly lack the precision necessary to recover meaningful trends.
Instead, we leverage recent advances in generative AI, which enable data extraction from unstructured text at scale. Our objective is to identify all records in the National Library of Medicine’s PubMed / MEDLINE database, from 2010 forward, that disclose the results of a clinical trial that studied the effects of a medicine in human subjects. We aim to classify records based only on the contents of the publication’s abstract, with an accuracy close to that of a human labeler. [^7] Many records in PubMed / MEDLINE have full-texts available for download, via the PubMed Central database. However, these full-texts are subject to copyright restrictions. Roughly one quarter of records in PubMed—$8$ million of the $34$ million total records—have full-texts that are available under licenses that would permit this type of research use (the PubMed Central Open Access Subset, at https://www.ncbi.nlm.nih.gov/pmc/tools/openftlist/). Thus, we restrict our focus to abstracts, for which no such restrictions apply.
We proceed in stages. Given the abstract of a scientific publication, we prompt two proprietary language models, OpenAI’s GPT-3.5 and GPT-4 (Ouyang et al.,, 2022; Achiam et al.,, 2023; Bubeck et al.,, 2023), to label whether the record satisfies our desired sample restrictions. We find considerable variation in performance with seemingly small differences in prompt text. We iteratively revise prompt text until small changes yield no improvement in performance. GPT-4 performs quite well off-the-shelf. The performance of GPT-3.5 is unsuitable for our application. However, use of GPT-4 at scale—to classify millions of records—is computationally and financially impractical. We classify a moderate number of abstracts with the best-performing prompts. The labels extracted from the proprietary language models are used as training data to fine-tune a set of open-source models. Our preferred sample is constructed using an ensemble of fine-tuned models, which matches the performance of the best proprietary model. False positive and false negative rates are below 5 percent. When this model is used to label the full sample of PubMed records, it yields a sample of approximately 150,000 publications that report the results of clinical trials from 2010 to 2022.
Three technical features of this fine-tuning process are worth highlighting. First, we document a phase-transition in model quality in the number of training labels used. Put differently, we observe a threshold in the number of training labels used in fine-tuning, beyond which the performance of the fine-tuned model increases dramatically. Second, models fine-tuned with labels extracted from GPT 3.5 exceed the performance of GPT 3.5. Third, models fine-tuned with labels extracted from GPT-4 match the performance of GPT-4 itself. That is, we show that researchers can distill the full benefits of large proprietary models into light-weight, resource-efficient models that can be realistically applied to standard data construction tasks.
@@
The resulting data allow us to document two sets of facts about the quantity of medical research over time. First, we find that the quantity of clinical trials has been effectively constant, year-on-year, across our period of interest. We collect a set of ex-ante and ex-post quality measures, derived from scientific publication data. On these dimensions, we find considerable heterogeneity that is, to our knowledge, new to the literature. Roughly half of clinical trial publications in our sample are never cited by a leading medical journal. Nearly 15 percent are never cited by any other scientific publication. However, we find that the distribution across clinical trials of measures of quality, importance, and geography is unchanging.
Second, our data allow us to shed light on the source of the classification errors that drive the trends documented in Bloom et al., (2020). [^8] These data are relevant outside of this particular context, for studies of medical innovation and for studies in the health services literature that use clinical trial publications to construct prescribing guidelines for physician practice. Identifying trial records in publication data has been a challenge for this health services literature, where existing best practices have both high true and false positive rates. See Thomas et al., (2021). As similar trends in the total quantity of scientific publications have been documented elsewhere, including in Park et al., (2023) and Chu and Evans, (2021), these insights may be more generally applicable to discussions about how large increases in the quantity of publications should be interpreted. We contrast trends in clinical trial production with trends in the production of medical research that is similar—conceptually and in its textual presentation in scientific publications—to clinical trials. In our preferred version, this set of other medical research is defined as the set of papers that cite clinical trials. [^9] In Section 4, we inspect the contents of this set and document that nearly all records fall into one of three categories: observational studies, preclinical studies, and meta-analyses/literature reviews. While the quantity of clinical trials is constant over this period, the quantity of this other non-trial research has increased by a factor of two.
Three facts about shifts in the composition and quality of these publications provide insight into the source of measured increases in research quantity. First, the large increase in the total number of non-trial papers in our sample is largely driven by researchers in China. Second, this rise in quantity coincides with a decrease (on available measures) in the average quality of publications. Third, there is a substantial increase in the quantity of meta-analyses and literature reviews. The main conclusion that emerges from this analysis is that errors in the classification of scientific publications can yield spurious findings about trends in research effort.
@@ 1.1. Related Literature
We draw on, and contribute to, a long literature on the measurement of invention. Decades of work have focused, in particular, on the development of proxies for the outputs of technological innovation. An especially prominent approach, due to Hall et al., (2000), uses citation-weighted counts of patents. Our focus, by contrast, is on the difficulties associated with measuring the inputs of innovation, which can similarly bias estimates of macroeconomic trends.
The new methods developed in this paper yield measures that call into question a fact at the center of debates involving health, innovation, and fiscal policy. Discussions of declining medical and pharmaceutical research productivity have been especially extensive in the popular press. Collison and Nielsen, (2018), Broad, (2023), Thompson, (2021), and Piper, (2023), for example, interpret trends in pharmaceutical research productivity, alongside patterns collected from other fields, as a “warning sign” for the future of technological progress (Piper,, 2023).
@@
A nascent literature investigates the use of large language models in economics (Dell,, 2024; Korinek,, 2023). We join a small set of papers—notably, Bartik et al., (2023) and Dell, (2024)—that use generative AI as a tool to construct high-quality economic data. Our approach is distinct, in that we fine-tune an open-source language model using data extracted from a larger, proprietary model. In doing so, we show that, with model distillation, it is feasible for researchers to achieve the quality of hand-constructed data at a practically interesting scale, with a reproducible and transparent methodology, at a feasible cost. Prominent examples of model distillation in the computer science literature focus on the construction of chat bots by fine tuning open source models with data extracted from proprietary models (Taori et al.,, 2023; Chiang et al.,, 2023; Xu et al.,, 2023). Other papers have fine-tuned models for simple completion tasks like arithmetic (Liu and Low,, 2023) or querying an API (Patil et al.,, 2023). We are distinguished from this literature by considering a problem with, arguably, a higher-degree of complexity: the construction of data of sufficient quality to be of substantive interest to social science.
@@
2. Identifying Clinical Trials

### v5

@@
We develop a method for assigning high-quality labels to unstructured text. This method is based on fine-tuning an efficient, open-source language model with data extracted from a large, proprietary language model. We apply this method to construct a census of published clinical trials. With these data, we revisit a literature that contends that pharmaceutical sector productivity is declining. Central to this conclusion are measurements of substantial increases in the quantity of clinical trials over time, unmatched by trends in measures of output. In our data, the quantity, quality, and composition of clinical trials are stable since 2010. We show that previous measurements are an artifact of biases introduced by shifts in the composition of other forms of research.
@@ 1. Introduction
For nearly thirty years, policy discussions about the pharmaceutical industry have been colored by the concern that sectoral productivity is in a state of decline (see e.g., Engelberg,, 1998; Harris,, 2011). A series of influential analyses, drawing on industry reports and U.S. Food and Drug Administration records, suggest that the resources required to bring a new drug to market have been increasing substantially year-on-year. Scannell et al., (2012), for example, estimate that the productivity of the pharmaceutical sector has halved roughly every decade since 1950; Ruffolo, (2006), Pammolli et al., (2011), Myers and Pauly, (2019), and Bloom et al., (2020) arrive at similar conclusions.
The policy consequences of these findings have been substantial. Proposals that would restructure public subsidies administered by the National Institutes for Health (NIH) (e.g., Executive Office of the President,, 2019; Moss et al.,, 2020; The Economist,, 2022), relax the regulatory standards for approval of new medicines by the U.S. Food and Drug Administration (FDA) (e.g., Roy,, 2021; Chertman,, 2023; Graboyes and Reimschisel,, 2017), and prevent the Centers for Medicare and Medicaid Services (CMS) from negotiating drug prices (e.g., York,, 2023; Long,, 2023; Inflation Reduction Act,, 2022) are premised, in no small part, on the concern that existing policy is contributing to this pattern of decline. [^1] In recent years, discussions of declining medical and pharmaceutical research productivity have been especially extensive in the popular press. Collison and Nielsen, (2018), Broad, (2023), Thompson, (2021), and Piper, (2023), for example, interpret trends in pharmaceutical research productivity, alongside patterns collected from other fields, as a ‘warning sign” for the future of technological progress (Piper,, 2023). For the same reasons, an architect of the 1984 Hatch Waxman Act, which established the modern system for pharmaceutical regulation in the United States, called for an overhaul of that regulatory system just fifteen years after its inception (Engelberg,, 1998). These patterns also spurred the 2011 establishment of the National Center for Translational Sciences, a federal agency with a roughly one billion dollar annual budget and a mandate to speed up the commercialization of life-saving medicines—on the grounds that, left to private firms, this process had become too inefficient. [^2] The fiscal year 2025 budget request for NCATS is \$926.1 million. For details on its founding, see Harris, (2011). For details on its budget, see https://ncats.nih.gov/about/budget.
Though this declining pharmaceutical sector productivity is often characterized as an empirical regularity on par with Moore’s Law in computing, its empirical support is quite thin. [^3] Moore’s Law is the observation that the number of transistors on an integrated circuit will double every two years with minimal rise in cost. See Moore, (1965). Scannell et al., (2012) coined “Eroom’s Law” (Moore, spelled backwards) to describe pharmaceutical sector productivity. On some measures, the quantity of resources devoted toward the commercialization of new medicines appears to be sharply increasing over time. Given that the number of innovative drugs approved by the FDA has been roughly constant, year-on-year, since 1990, this divergence between inputs and outputs generates the conclusion of declining productivity.
This paper constructs new data on pharmaceutical sector inputs. We develop a method for assigning high-quality labels to unstructured text. This method is based on fine-tuning an efficient, open-source language model with data extracted from a large, proprietary language model. We apply this method to construct a new census of clinical trials disclosed in scientific publications. We document that the quantity, quality, and composition of published clinical trials have remained remarkably stable year-over-year. On the basis of these facts, we find that there is insufficient evidence to reject the null hypothesis that pharmaceutical sector productivity has remained stable since 2010.
We show that reported increases in the quantity of published clinical trials, offered as recent evidence of declining productivity in Bloom et al., (2020), result from imprecise classification methods. These methods misattribute shifts in the composition of biomedical research to trends in clinical trial production. When our data are combined with evidence from Sertkaya et al., (2024)—that clinical trial costs, and broader drug development expenses, have remained roughly constant since 2000—our findings leave little in the way of evidence for the conventional wisdom of a persistent decline in pharmaceutical sector productivity.
This paper makes two primary contributions. First, we revise a fact central to debates on biomedical productivity. Second, we demonstrate how advances in generative AI can address longstanding challenges in measuring innovation. The paper is, thus, organized in two parts.
@@
In the first part, we introduce a procedure for constructing task-specific large language models, particularly suited to data construction problems that involve classification. [^4] We include, as a supplemental appendix to this paper, a more detailed summary of this procedure, which may be useful for future work. The class of methods that we consider has two steps. First, a comparatively small number of labels are extracted from a large, typically proprietary, language model. [^5] By proprietary large language model, we are referring, at present, to models such as OpenAI’s GPT 3.5 and GPT-4. Alternative models include Anthropic’s Claude 3.5 Opus, Google’s Gemini, among many others. Second, these labels are used as training data for a smaller, more efficient, open-source language model. [^6] Here, we are referring to classes of open-source language models such as RoBERTa, based on Google’s BERT model, among others. This process is an instance of model distillation, a growing area of methodological research in computer science (Taori et al.,, 2023; Xu et al.,, 2024). This procedure allows researchers to capture the benefits of complex, frontier models, while retaining the resource efficiency and transparency of lighter-weight open-source models.
We apply this method in stages. Our objective is to identify all records in the National Library of Medicine’s PubMed/MEDLINE database, from 2010 forward, that disclose the results of a clinical trial that studied the effects of a medicine in human subjects. We aim to classify records based only on the contents of the publication’s abstract, with an accuracy close to that of a human labeler. We devise an interface to iteratively revise model prompts. With the highest performing prompts, we use proprietary language models, OpenAI’s GPT-3.5 and GPT-4, to classify a large number of records (Achiam et al.,, 2023; Bubeck et al.,, 2023). These noisy labels are used as training data to fine-tune a set of open-source models. An ensemble of the fine-tuned models matches the performance of the best proprietary model, with false positive and false negative rates below 5 percent, and produces a sample of approximately 150,000 publications that report the results of clinical trials from 2010 to 2022.
@@
The second part of this paper interprets trends in these data, in light of existing evidence. We begin by establishing that trends in clinical trials, as measured in publication data, are central to conclusions of declining biomedical and pharmaceutical productivity. We show that standard methods for data classification identify clinical trials in publication data imprecisely. These imprecise measurements indicate, spuriously, that the quantity of clinical trials has increased sharply since 2010. With improved classification, we find that the quantity, quality, and composition of clinical trials have been remarkably stable over this period.
@@
We collect a set of ex-ante and ex-post quality measures, derived from scientific publication data. On these dimensions, we find considerable heterogeneity that is, to our knowledge, new to the literature. Roughly half of clinical trial publications in our sample are never cited by a leading medical journal. Nearly 15 percent are never cited by any other scientific publication. However, we find that the distribution across clinical trials of measures of quality, importance, and geography is unchanging over time.
Our data offer more suggestive evidence on a second piece of evidence cited in support of declining research productivity—that the quantity of scientific publications is increasing sharply over time. We document evidence of large compositional shifts in other biomedical publications, including a shift in the geographic composition of research and a large increase in the number of articles that summarize existing literature. In line with the arguments of a long-standing literature on similar shifts in patent data (e.g., Griliches,, 1990, 1994), these patterns cast doubt on whether simple counts of scientific publications are stable, informative measures of research effort. These insights may be more generally applicable to discussions about how large increases in the quantity of publications—as documented in Park et al., (2023) and Chu and Evans, (2021)—ought to be interpreted.
@@ 1.1. Related Literature
We draw on, and contribute to, a long literature on the measurement of invention. This paper considers, in effect, a binary classification task, of a kind that has long been a challenge for studies of innovation. The proliferation of large-scale corpora of patents and scientific publications has, seemingly, expanded opportunities to measure instances of research and invention (Bryan and Williams,, 2021). Patents and publications are attractive proxies for innovation, as they necessarily relate to instances of invention and research. Their utility is limited, however, by heterogeneity in their technical content and value (see e.g., Pakes,, 1986; Griliches,, 1990, 1994). Consequently, trends in counts of documents are often confounded by trends in their composition (Lerner and Seru,, 2022). Large-scale classification efforts have focused on mapping pieces of unstructured text to economic quantities of interest. [^7] The most prominent example of such a classification effort is the construction of the National Bureau of Economic Research (NBER) Patent Data, which, among other things, maps patent records to technology classes (Hall et al.,, 2001). Earlier, Scherer, (1984) classified, by hand, over 15,000 patents according to their technological category. In this paper, we demonstrate that large language models can facilitate classification tasks of this kind, achieving an accuracy and precision comparable to hand-made labels at a research-relevant scale and budget.
More specifically, within studies of medical innovation, we contribute to a body of work that uses clinical trials as a measure of research investment. Trials—discrete, regulated experiments that generate structured statistical information on new and existing health technologies (Chavez-MacGregor and Giordano,, 2016)—are an especially attractive proxy for the level and composition of medical research. Since the early 2000s, clinical trials have been reported in a consistent, standardized format across scientific publications and, with less consistency, in administrative databases (DeVito et al.,, 2020; DeVito and Goldacre,, 2021). Trials have, thus, been widely used to measure relative and absolute differences in the composition of research investment (see e.g., Finkelstein,, 2004; Yin,, 2008; Budish et al.,, 2015). Recent studies of information production and disclosure have examined the substantive content of trials (see e.g., Alsan et al., (2024), Oostrom, (2024), and Kao and Oostrom, (2024)).
Given limitations in publicly-available administrative trial registries (Zarin et al.,, 2017; DeVito et al.,, 2020; DeVito and Goldacre,, 2021), recent empirical studies in economics have used trial data extracted from scientific publications (e.g., Lichtenberg,, 2018; Bloom et al.,, 2020; Kao,, 2024; Guzman et al.,, 2024). The challenge, for each, has been again one of classification: of the millions of records in large-scale publication databases, which disclose the results of a clinical trial? Manually identifying clinical trial records at a scale sufficient to study aggregate trends is infeasible. [^8] A small number of papers curate select samples of clinical trial data, for specific research questions. For example, Oostrom, (2024) collects records of psychiatric clinical trials as identified in prominent meta-analyses for a study of the relationship between sources of clinical trial funding and reported drug efficacy. We show that standard methods for classification—based on publication data and machine learning tools—perform poorly. In particular, we show that keyword-based searches, including methods used in a productivity analysis in Bloom et al., (2020), are very imprecise and generate spurious measurements of increases in the quantity of clinical trials. [^9] Specifically, Bloom et al., (2020) categorize many records as clinical trials that discuss trials (e.g., meta-analyses and literature reviews). We show, in Section 4, that changes in the scientific publishing landscape increase the size of this categorization error over time. Specialized classifiers developed for use in the health services literature, as in Thomas et al., (2021), similarly lack the precision necessary to recover meaningful trends. We introduce both a new method and new data that address this problem.
@@
A nascent literature investigates the use of large language models in economics (Dell,, 2024; Korinek,, 2023; Ludwig et al.,, 2025). We join a small set of papers—notably, Bartik et al., (2023) and Dell, (2024)—that use generative AI as a tool to construct high-quality economic data. Our approach is distinct, in that we fine-tune an open-source language model using data extracted from a larger, proprietary model. In doing so, we show that, with model distillation, it is feasible for researchers to achieve the quality of hand-constructed data at a practically interesting scale, with a reproducible and transparent methodology, at a feasible cost. Prominent examples of model distillation in the computer science literature focus on the construction of chat bots by fine tuning open source models with data extracted from proprietary models (Taori et al.,, 2023; Chiang et al.,, 2023; Xu et al.,, 2023). Other papers have fine-tuned models for simple completion tasks like arithmetic (Liu and Low,, 2023) or querying an API (Patil et al.,, 2023). We are distinguished from this literature by considering a problem with, arguably, a higher-degree of complexity: the construction of data of sufficient quality to be of substantive interest to social science.
@@
2. Counting Clinical Trials